# Stacked conformal prediction

### v1

@@ 4 Specializing the meta-learner
Using multiple linear regression as the meta-learner of the feasible stack, we implemented the stacked conformal prediction procedure for two examples. The California housing dataset (Pace and Barry, 1997) comprises information on 20,640 census tracts in California, drawn from the 1990 U.S. Census. It includes eight predictors and the response variable is the median house value in US dollars for each tract. The Ames housing dataset (De Cock, 2011) contains information on 2,930 houses sold in Ames, Iowa from 2006 to 2010. For the Ames dataset, we have eighty predictors and the response variable is the sale price of the house in US dollars.
@@
We used Random Forests (Breiman, 2001) and CatBoost (Prokhorenkova et al., 2018) as the base-learners of the stack. For both datasets, we used 70% of the available data as our training sample, and generated prediction sets for the remaining 30% test sample units. For comparison, we used conformalized quantile regression (CQR) (Romano et al., 2019) as an inductive alternative. The Quantile Random Forest algorithm (Meinshausen, 2006) was used in the implementation of CQR. For CQR, the calibration samples sizes were chosen to be 2,000 and 500, for the California and Ames datasets, respectively. In both examples, we used a 90% nominal miscoverage level. Figures 1 and 2 show the prediction intervals for fifty randomly chosen test sample units. Table 1 summarizes the results, showing that the stacked conformal prediction intervals exhibited appropriate empirical coverage, with shorter median length, when compared to the CQR results.
@@
Figure 1: Prediction intervals for fifty test sample units in the California Housing dataset.
@@
Figure 2: Prediction intervals for fifty test sample units in the Ames Housing dataset.
@@
Table 1: Prediction interval length summaries (US dollars) and empirical coverage.
Dataset Method 1st quartile Median 2nd quartile Coverage
California Stacked CP 96,927 119,003 147,988 89.9%
CQR 104,433 156,600 231,500 90.8%
Ames Stacked CP 41,652 59,478 88,088 91.1%
CQR 44,530 68,960 108,839 90.2%
5 Open source implementations
R (R Core Team, 2024) and Python (Python Core Team, 2023) codes for the full reproduction of the examples in the paper are available at:

### v2

@@ 18 $r_{0}\leftarrow y^{(\text{res})}_{0}/(1+\hat{\delta}_{0})$
We implemented the stacked conformal prediction procedure for two examples. The California housing dataset (Pace and Barry, 1997) comprises information on 20,640 census tracts in California, drawn from the 1990 U.S. Census. It includes eight predictors and the response variable is the median house value in US dollars for each tract. The Ames housing dataset (De Cock, 2011) contains information on 2,930 houses sold in Ames, Iowa from 2006 to 2010. For the Ames dataset, we have eighty predictors and the response variable is the sale price of the house in US dollars.
@@
We used Random Forests (Breiman, 2001) and CatBoost (Prokhorenkova et al., 2018) as the stack base-learners. For both datasets, we used 70% of the available data as our training sample, and generated prediction sets for the remaining 30% test sample units. For comparison, we used conformalized quantile regression (CQR) (Romano et al., 2019) as an inductive alternative. The Quantile Random Forest algorithm (Meinshausen, 2006) was used in the implementation of CQR. For CQR, the calibration samples sizes were chosen to be 2,000 and 500, for the California and Ames datasets, respectively, following the calibration sample size discussion in Marques F. (2025). For each dataset, Figures 1 and 2 show the prediction intervals for fifty randomly chosen test sample units, using a 90% nominal coverage level ($\alpha=0.1$). Table 1 summarizes the results for different nominal coverage levels, showing that the stacked conformal prediction intervals exhibited appropriate empirical coverage, with shorter median interval width, when compared to the CQR results.
@@
Figure 1: Prediction intervals for fifty test sample units in the California Housing dataset, using a 90% nominal coverage level ($\alpha=0.1$).
@@
Figure 2: Prediction intervals for fifty test sample units in the Ames Housing dataset, using a 90% nominal coverage level ($\alpha=0.1$).
@@
Table 1: Empirical coverage and prediction interval width summaries (US dollars) for different nominal miscoverage levels $\alpha$.
Dataset Method $1-\alpha$ Empirical coverage 1st quartile Median 3rd quartile
California Stacked CP 80% 80.9% 69,226 84,995 105,696
California CQR 80% 80.1% 67,485 105,580 162,900
California Stacked CP 85% 85.5% 80,864 99,283 123,470
California CQR 85% 85.7% 81,111 125,310 192,075
California Stacked CP 90% 89.9% 96,927 119,003 147,988
California CQR 90% 90.8% 104,433 156,600 231,500
Ames Stacked CP 80% 81.9% 27,665 39,511 58,467
Ames CQR 80% 81.2% 29,796 48,355 79,813
Ames Stacked CP 85% 86.7% 32,645 46,624 69,001
Ames CQR 85% 85.9% 34,401 55,210 91,590
Ames Stacked CP 90% 91.1% 41,652 59,478 88,088
Ames CQR 90% 90.2% 44,530 68,960 108,839
5 Concluding remarks and open source implementations
We have presented the stacked conformal prediction procedure in a regression setting, but its general structure readily extends to other tasks – classification being a prime example. The key requirement in any such extension is the choice of a meta-learner that admits full conformalization at manageable computational cost. In classification problems, a natural meta-learner candidate would be the $k$-nearest neighbors classifier, equipped with a distance metric tailored to aggregate the class-probability outputs produced by the stack base-learners. R (R Core Team, 2024) and Python (Python Core Team, 2023) codes, including C++ (Stroustrup, 2013) extensions, for full reproduction of the examples in the paper are available at:
@@ 5 Concluding remarks and open source implementations
Appendix: Proofs
Proof .1.
Laying out the second-level pairs as the $(n+1)\times(M+1)$ matrix
$U=g(T,Q)$ $=\begin{pmatrix}Z_{1}&Y_{1}\\ Z_{2}&Y_{2}\\ \vdots&\vdots\\ Z_{n}&Y_{n}\\ Z_{n+1}&Y_{n+1}\end{pmatrix}=\begin{pmatrix}Z_{1,1}&Z_{1,2}&\cdots&Z_{1,M}&Y_{1}\\ Z_{2,1}&Z_{2,2}&\cdots&Z_{2,M}&Y_{2}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ Z_{n,1}&Z_{n,2}&\cdots&Z_{n,M}&Y_{n}\\ Z_{n+1,1}&Z_{n+1,2}&\cdots&Z_{n+1,M}&Y_{n+1}\end{pmatrix}$
$=\begin{pmatrix}\hat{\mu}_{1}\!\left(Q_{\setminus\phi(1)}T,X_{1}\right)&\hat{\mu}_{2}\!\left(Q_{\setminus\phi(1)}T,X_{1}\right)&\cdots&\hat{\mu}_{M}\!\left(Q_{\setminus\phi(1)}T,X_{1}\right)&Y_{1}\\ \hat{\mu}_{1}\!\left(Q_{\setminus\phi(2)}T,X_{2}\right)&\hat{\mu}_{2}\!\left(Q_{\setminus\phi(2)}T,X_{2}\right)&\cdots&\hat{\mu}_{M}\!\left(Q_{\setminus\phi(2)}T,X_{2}\right)&Y_{2}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ \hat{\mu}_{1}\!\left(Q_{\setminus\phi(n)}T,X_{n}\right)&\hat{\mu}_{2}\!\left(Q_{\setminus\phi(n)}T,X_{n}\right)&\cdots&\hat{\mu}_{M}\!\left(Q_{\setminus\phi(n)}T,X_{n}\right)&Y_{n}\\ \hat{\mu}_{1}\!\left(Q_{\setminus\phi(n+1)}T,X_{n+1}\right)&\hat{\mu}_{2}\!\left(Q_{\setminus\phi(n+1)}T,X_{n+1}\right)&\cdots&\hat{\mu}_{M}\!\left(Q_{\setminus\phi(n+1)}T,X_{n+1}\right)&Y_{n+1}\end{pmatrix},$
choose any permutation matrix $\Pi\in\mathbb{S}_{n+1}$, and consider the permutation of the $U$ rows defined by $\Pi U=\Pi g(T,Q)$. This permuted matrix $\Pi g(T,Q)$ can be obtained in a second way. Consider the folding scheme matrix defined by $Q\Pi^{-1}$. Because $(Q\Pi^{-1})(\Pi T)=QT$, this folding scheme matrix $Q\Pi^{-1}$ has the special property that it assigns the training sample units in the permuted training sample $\Pi T$ to the exact same folds that they had been assigned by the folding scheme matrix $Q$ applied to the original training sample $T$. Therefore, the assumed training data symmetry (italic-$\star$2) of the base learners $\hat{\mu}_{m}$ in the definition of the symmetric stack yields the property: ${\Pi g(T,Q)=g(\Pi T,Q\Pi^{-1})}$. Since the class of permutation matrices is closed under multiplication and taking inverses, we have, for every $q\in\mathbb{S}_{n+1}$, that $\Pr(Q\Pi^{-1}=q)=\Pr(Q=q\Pi)=1/(n+1)!$, implying that $Q$ and $Q\Pi^{-1}$ have the same distribution. Because we assumed that the training sample $T$ is exchangeable, and that $T$ and the folding scheme matrix $Q$ are independent, we come to the conclusion that $(T,Q)$ and $(\Pi T,Q\Pi^{-1})$ have the same distribution. Hence, $U=g(T,Q)$ and $\Pi U=\Pi g(T,Q)=g(\Pi T,Q\Pi^{-1})$ have the same distribution, as desired.
The following result is used in the proof of Proposition 2.2.
Lemma .2.
Let $U_{1},U_{2},\dots,U_{n}$ be exchangeable random variables, and denote their order statistics by ${U_{(1)}\leq U_{(2)}\leq\dots\leq U_{(n)}}$. It follows that $P(U_{i}\leq U_{(k)})\geq k/n$, for each ${i,k=1,\dots,n}$.
Proof .3.
Considering that we may have ties, it follows from the definition of $U_{(k)}$ that at least $k$ of the $U_{i}$’s are less than or equal to $U_{(k)}$. Hence, ${\sum_{i=1}^{n}I_{\{U_{i}\leq U_{(k)}\}}\geq k}$, almost surely. Taking expectations and noting that, by exchangeability, the probability $P(U_{i}\leq U_{(k)})$ is the same for every ${i=1,\dots,n}$, the result follows.
Proof .4.
Taking $y=Y_{n+1}$, the definitions in the proposition statement and the result in Proposition 2.1 imply that the random conformity scores
$R_{1}^{(Y_{n+1})},R_{2}^{(Y_{n+1})},\dots,R_{n}^{(Y_{n+1})},R_{n+1}^{(Y_{n+1})}$
are exchangeable. Using Lemma .2 above, we have that
$\Pr\!\left(R^{(Y_{n+1})}_{n+1}\leq R^{(Y_{n+1})}_{(k)}\right)\geq\frac{k}{n+1},$
for $k=1,\dots,n+1$. Choosing $k=\lceil(1-\alpha)(n+1)\rceil\geq(1-\alpha)(n+1)$, the result follows, since the definition of the random prediction set $C_{n+1}^{(\alpha)}$ entails that $Y_{n+1}\in C_{n+1}^{(\alpha)}$ if and only if $R^{(Y_{n+1})}_{n+1}\leq R^{(Y_{n+1})}_{(\lceil(1-\alpha)(n+1)\rceil)}$.
Proof .5.
Introduce the notations
$r=R_{(\lceil(1-\alpha)(n+1)\rceil)}^{(Y_{n+1})}\quad\text{and}\quad\tilde{r}=\tilde{R}_{(\lceil(1-\alpha)(n+1)\rceil)}^{(Y_{n+1})},$
and define the event
$B_{\epsilon}=\left\{\max\left\{\left|\tilde{R}_{n+1}^{(Y_{n+1})}-R_{n+1}^{(Y_{n+1})}\right|,\left|\tilde{r}-r\right|\right\}<\epsilon/2\right\}.$
If the statement in $B_{\epsilon}$ is true and $R_{n+1}^{(Y_{n+1})}\leq r-\epsilon$, then $\tilde{R}_{n+1}^{(Y_{n+1})}<R_{n+1}^{(Y_{n+1})}+\epsilon/2\leq r-\epsilon/2<\tilde{r}$.
Therefore, using the lower bound in Proposition 2.2, the result follows:
$\Pr\!\left(Y_{n+1}\in\tilde{C}_{n+1}^{(\alpha)}\right)$ $=\Pr\!\left(\tilde{R}_{n+1}^{(Y_{n+1})}\leq\tilde{r}\right)$
$\geq\Pr\!\left(\left\{\tilde{R}_{n+1}^{(Y_{n+1})}<\tilde{r}\right\}\cap B_{\epsilon}\right)$
$\geq\Pr\!\left(\left\{R_{n+1}^{(Y_{n+1})}\leq r-\epsilon\right\}\cap B_{\epsilon}\right)$
$=\Pr\!\left(R_{n+1}^{(Y_{n+1})}\leq r-\epsilon\right)-\Pr\!\left(\left\{R_{n+1}^{(Y_{n+1})}\leq r-\epsilon\right\}\cap B^{c}_{\epsilon}\right)$
$\geq\Pr\!\left(R_{n+1}^{(Y_{n+1})}\leq r-\epsilon\right)-\Pr\!\left(B^{c}_{\epsilon}\right)$
$\geq\Pr\!\left(R_{n+1}^{(Y_{n+1})}\leq r-\epsilon\right)-\delta$
$=\Pr\!\left(R_{n+1}^{(Y_{n+1})}\leq r\right)-\Pr\!\left(r-\epsilon<R_{n+1}^{(Y_{n+1})}\leq r\right)-\delta$
$=\Pr\!\left(Y_{n+1}\in C_{n+1}^{(\alpha)}\right)-\Pr\!\left(r-\epsilon<R_{n+1}^{(Y_{n+1})}\leq r\right)-\delta$
$\geq 1-\alpha-\delta-\Pr\!\left(r-\epsilon<R_{n+1}^{(Y_{n+1})}\leq r\right)$
$=1-\alpha-\delta-h(\epsilon).$